OpenStudy (amistre64):

Every odd integer is the difference of 2 squares. Direct Proof. Let m and n be any integers such that; 2m^2 – (2n+1)^2 = 2k +-1 4m^2 – 4n^2 -4n -1 = 2k +- 1 2(2m^2 -2n^2 -2n) – 1 = 2k +- 1 Does this work?

OpenStudy (amistre64):

(2m)^2 – (2n)^2 = 2k +-1 4m^2 – 4n^2 = 2k +- 1 2(m^2 -n^2) = 2k +- 1 Doesnt seem to fit

OpenStudy (amistre64):

every odd integer is the difference of 2 consecutive squares

myininaya (myininaya):

do you mean 2 consecutive squares or do you mean the square of 2 consecutive numbers ?

OpenStudy (amistre64):

1 - 0 = 1 2 - 1 = 3 3 - 2 = 5 4 - 3 = 7 5 - 4 = 9.

OpenStudy (amistre64):

thoseare lazy squares :)

myininaya (myininaya):

3-2=5?

OpenStudy (amistre64):

3^2 - 2^2 = 5 :)

myininaya (myininaya):

oh lazy squares i get it

myininaya (myininaya):

ok oh thats neat never realized that i don't think maybe i forgot let n be the fist integer and let n+1 be the next integer (n+1)^2-n^2=n^2+2n+1-n^2=2n+1 2n+1 is odd

myininaya (myininaya):

so (n+1)^2-n^2 is odd

myininaya (myininaya):

go ahead satellite rain on my parade

OpenStudy (anonymous):

your last line is the only one needed right?

myininaya (myininaya):

yes

OpenStudy (amistre64):

yes

myininaya (myininaya):

maybe

OpenStudy (amistre64):

i was wondering if the proof had to be that every odd number is the difference of any 2 squares; but i might be reading to much into it

OpenStudy (anonymous):

of course. general odd integer \[2k+1=(k+1)^2-k^2\]

OpenStudy (amistre64):

and of course the teacher said that k would indicate the same integer

OpenStudy (amistre64):

would have to include subs

OpenStudy (anonymous):

as in the proof that \[\sum_{k=0}^n 2k+1=n^2\]

OpenStudy (anonymous):

start with \[\sum_{k=0}^n(k+1)^2-k^2=n^2\] and then do the algebra to see that you get \[\sum_{k=0}^n 2k+1\]

OpenStudy (amistre64):

we havent learnted no sum notation so its not allowed ... lol

OpenStudy (anonymous):

lol. what you wrote above is exactly the same thing right? like the telescoping sum. one term minus the previous one

OpenStudy (amistre64):

it is... but I gots the antiGuass teacher ...

OpenStudy (amistre64):

that aint how I showed it to you !!! YOU FAIL!!!

OpenStudy (anonymous):

antiguasse? lol

OpenStudy (anonymous):

think photoshop comes with an anti gausse feature

OpenStudy (amistre64):

i believe it does, but you have to read it in korean